# Hao Ma1, Zhiyuan Peng2, Xu Li3, Yukai Li1, Mingjie Shao1, Qiuqiang Kong4, and Ju Liu1

### v1

@@ III-C Results and Analysis
Language-free versus weakly-supervised. We conduct weakly-supervised experiments by invoking the label-formed annotations on AudioSet and transforming the audio labels into captions by prefixing each with “The sound of ”. The detailed results on each of the evaluation benchmarks are shown in Table I. Comparing the weakly-supervised training scheme with our proposed language-free training scheme, the weakly-supervised method suppresses the language-free methods only on AudioSet but falls behind on other evaluation benchmarks. This is because the query language in the AudioSet test set is defined in the same way as in weakly-supervised training. This phenomenon shows that our proposed approach has better generalizability to various query forms, especially those in natural language form, as demonstrated by the results on AudioCaps and Clotho v2.
@@
Ablation study of Gaussian perturbation strength. The choice of the Gaussian perturbation variance has a significant effect on the performance of language-free training. To determine the Gaussian perturbation strength, we vary the variation of injected Gaussian noise from 0 to 1e-1 and train a TSE model for each value. We run evaluations on all five testing datasets and present the results of averaged SDRi and SI-SDRi in Fig. 2.
@@
The green lines in Fig. 2 illustrate the effect of noise variance $\epsilon$ on the performance of the standard Gaussian noise injection strategy, which injects Gaussian noise directly into the target audio embedding to form the condition embedding in training. As shown in Fig. 2, the appropriate choice of noise variance has a crucial impact on the effectiveness of this approach. Specifically, when the noise variance is small, the modality gap and information leakage will cause a significant drop in performance. On the other hand, if the noise variance is too large, the useful information in audio embeddings is overwhelmed by too much noise. According to Fig. 2, we set the noise variance to 1e-2 for all subsequent experiments for the standard Gaussian noise injection strategy.
@@
The red lines in Fig. 2 illustrate the effect of the noise variance $\epsilon_{1}$ on the performance of our proposed retrieval-augmented strategy. The experimental results show that for our proposed retrieval-augmented strategy, Gaussian noise injection of appropriate intensity also improves performance to some extent, but the effect is far less dramatic than for the standard noise injection strategy, further demonstrating that the better performance achieved by the proposed retrieval-augmented strategy does not depend on the noise perturbation.
@@
Ablation study of LLM-produced captions. Our proposed retrieval-augmented language-free training scheme requires retrieving corresponding text embeddings from a pre-encoded text embedding cache to construct condition embeddings. Therefore, the quality of this embedding cache plays a crucial role in the performance of the proposed retrieval augmented strategy. We suggest using LLM-produced audio captions to construct this embedding cache. In this section, we design experiments to verify the impact of LLM-produced audio captions on the performance of our proposed method. Specifically, we first establish a baseline system where we take all the 527 audio labels from AudioSet and construct 527 audio captions by prefixing each with “The sound of ”. These captions are then encoded into text embeddings using the CLAP text encoder and stored in an embedding cache, thus forming our baseline system. Then we invoke WavCaps [19] to simulate LLM-produced audio captions. We encode all the 403,050 LLM-produced audio captions from WavCaps into text embeddings, which are additionally added to the embedding cache in the baseline system. We train a new TSE model for each of the two settings and perform evaluation experiments on all five datasets. The results are shown in Fig. 3. The experimental results show that the proposed method, combined with the LLM-produced audio captions, achieves significant performance improvements on most evaluation benchmarks.
@@
Scale up the training data. In this section, we scale up the training data to include all collected audio clips from AudioSet and FreeSound to further validate the effectiveness of our method in leveraging large-scale non-parallel data. The corresponding results are presented in the last two rows of Table I. From these results, we can see that the proposed language-free training scheme performs better on all evaluation benchmarks after scaling up the training data. These results outperform the current SOTA supervised training schemes on most of the evaluated benchmarks, thus demonstrating the effectiveness of the proposed method in efficiently leveraging a large amount of audio data without annotation to train a language-queried target sound extraction model with strong generalizability.
Figure 3: Ablation study of LLM-produced captions.
@@ IV Conclusion
In this work, we present a novel language-free training paradigm for language-queried target sound extraction, leveraging the modality-aligned embedding space generated by the pre-trained CLAP model. By harnessing unlabeled audio clips, our model deftly sidesteps the necessity for labor-intensive text-audio pairings. We address challenges like modality gap and information leakage through a retrieval-augmented strategy, employing an embedding cache constructed from LLM-produced audio captions. Our extensive evaluation on several benchmarks demonstrates that this language-free training approach achieves consistent and notable performance improvements over existing state-of-the-art with better generalizability. This method significantly reduces the reliance on annotated data, offering a more adaptable and scalable solution for TSE.

### v3

@@ III-C2 Parallel-Data-Free Versus Weakly-Supervised
We conduct weakly-supervised experiments by invoking the label-formed annotations on AudioSet and transforming the audio labels annotated to each of the audio clips into captions by prefixing them with “The sound of ”. The detailed results on each of the evaluation benchmarks are shown in Table I. Comparing the weakly-supervised training scheme with our proposed parallel-data-free training schemes, the weakly-supervised method suppresses the parallel-data-free methods only on AudioSet but falls behind on other evaluation benchmarks. This is because the query language in the AudioSet test set is defined in the same way as in weakly-supervised training. This phenomenon shows that even when one-to-one weak labels (like those in AudioSet) are available, our method can achieve superior performance and better generalizability on natural language-formed user queries without requiring such annotations.
III-C3 Effect of Gaussian Perturbation Strength
@@
The choice of the Gaussian perturbation variance has a significant effect on the performance of parallel-data-free training [20, 25, 22]. To determine the Gaussian perturbation strength, we vary the variation of injected Gaussian noise from 0 to 1e-1 and train a TSE model for each value. We run experiments on a held-out validation set consisting of audio-text pairs from AudioCaps and Clotho v2 validation split and present the SDRi and SI-SDRi in Fig. 2.
The green lines in Fig. 2 illustrate the effect of noise variance $\epsilon$ on the performance of the vanilla strategy, which injects Gaussian noise directly into the target audio embedding to form the condition embedding in training. As shown in Fig. 2, the appropriate choice of noise variance has a crucial impact on the effectiveness of this approach. Specifically, when the noise variance is small, the mismatch between training and testing will cause a significant drop in performance. On the other hand, if the noise variance is too large, the useful information in audio embeddings is overwhelmed by too much noise. According to Fig. 2, we set the noise variance to 1e-2 for all subsequent experiments for the vanilla strategy.
@@
The red lines in Fig. 2 illustrate the effect of the noise variance $\epsilon_{1}$ on the performance of our proposed retrieval-augmented strategy. The experimental results show that for our proposed retrieval-augmented strategy, Gaussian noise injection of appropriate intensity also improves performance to some extent, but the effect is far less dramatic than that for the vanilla strategy, further demonstrating that the better performance achieved by the proposed retrieval-augmented strategy does not depend on the noise perturbation since the training-testing mismatch has already been effectively addressed by retrieving pre-encoded text embeddings.
III-C4 Effect of LLM-Produced Captions
@@
Our proposed retrieval-augmented parallel-data-free training scheme requires retrieving corresponding text embeddings from a pre-encoded text embedding cache to construct condition embeddings. Therefore, the quality of this embedding cache plays a crucial role in the performance of the proposed retrieval strategy. We suggest using LLM-produced audio captions to construct this embedding cache. In this section, we design experiments to verify the impact of LLM-produced audio captions on the performance of our proposed method. Specifically, we first establish a baseline system where we take all the 527 audio labels from AudioSet and construct 527 audio captions by prefixing each with “The sound of ”. These captions are then encoded into text embeddings using the CLAP text encoder and stored in an embedding cache, thus forming our baseline system. Then we invoke WavCaps [27] to simulate LLM-produced audio captions. We encode all the 403,050 LLM-produced audio captions from WavCaps into text embeddings, which are additionally added to the embedding cache in the baseline system. We train a new TSE model for each of the two settings and perform evaluation experiments on all five datasets. The results are shown in Fig. 3. The experimental results show that the proposed method, combined with the LLM-produced audio captions, achieves significant performance improvements on most evaluation benchmarks.
III-C5 Scale up the Training Data
@@
In this section, we scale up the training data to include all collected audio clips from AudioSet and FreeSound to further validate the effectiveness of our method in leveraging large-scale non-parallel data. The corresponding results are presented in the last two rows of Table I. From these results, we can see that the proposed parallel-data-free training schemes perform better on all evaluation benchmarks after scaling up the training data. These results also outperform the current SOTA-supervised training schemes on most of the evaluated benchmarks, thus demonstrating the effectiveness of the proposed method in efficiently leveraging a large amount of audio data without annotations to train a language-queried target sound extraction model with strong generalizability.
Figure 3: Effect of LLM-produced captions.
@@ IV Conclusion
In this work, we present a novel parallel-data-free training paradigm for language-queried target sound extraction, leveraging the modality-aligned embedding space generated by the pre-trained CLAP model. By harnessing unlabeled audio clips, our model deftly sidesteps the necessity for labor-intensive text-audio pairings. We address challenges like modality gap and information leakage through a retrieval-augmented strategy, employing an embedding cache constructed from LLM-produced audio captions. Our extensive evaluation on several benchmarks demonstrates that this parallel-data-free training approach achieves consistent and notable performance improvements over existing state-of-the-art with better generalizability. This method significantly reduces the reliance on human-annotated audio-text data pairs, offering a more adaptable and scalable solution for TSE.